OpenStudy (aravindg):

.If alpha and beta be 2 distinct roots satisfying equation a cos theta+bsin theta=c,Show that cos (alpha+beta)=(a^2-b^2)/(a^2+b^2)

OpenStudy (aravindg):

.............

OpenStudy (anonymous):

\[a\cos(\theta)+b\sin(\theta)=c\] prove that \[\cos(\alpha+\beta)=\frac{a^2-b^2}{a^2+b^2}\]

OpenStudy (aravindg):

ya

OpenStudy (anonymous):

OKay...I am posting the solution ..I got it

OpenStudy (anonymous):

\[\cos(\alpha+\beta) = \cos(\alpha)\cos(\beta)-\sin(\alpha)\sin(\beta)\]

OpenStudy (anonymous):

\[a\cos(\alpha)+b\sin(\alpha) = a\cos(\beta)+b\sin(\beta)\]

OpenStudy (anonymous):

\[cos(\alpha+\beta)= cos\alpha*cos\beta - sin\alpha*sin\beta\] \[acos\theta + bsin\theta=c\] \[acos\theta = c - bsin\theta\] \[\text{Square both sides}\] \[a^2 (1 - sin^\theta) = c^2 + b^2sin^2\theta -2*c*b*sin\theta\] \[(b^2 + a^2 )sin^2\theta -2 *c*b*sin\theta + c^2 -a^2=0\] \[\text{Product of Roots }= \frac{c}{a}\] \[sin\alpha*sin\beta=\frac{c^2-a^2}{a^2+b^2}\] \[\text{Now use similar method to have an equation in terms of } cos^2\theta\] \[acos\theta + bsin\theta=c\] \[acos\theta-c = -bsin\theta\] \[a^2cos^2\theta + c^2 - 2*a*ccos\theta = b^2 (1-cos^2\theta)\] \[(a^2 + b^2)cos^2\theta - 2*a*cos\theta*c +c^2 -b^2=0\] \[\text{Product of Roots }= \frac{c}{a}\] \[cos\alpha*cos\beta = \frac{c^2-b^2}{a^2+b^2}\] \[\text {Now substitute the values in First Equation }\] \[cos(\alpha+\beta)= cos\alpha*cos\beta - sin\alpha*sin\beta\] \[\implies \frac{c^2 -b^2 + a^2 -c^2}{a^2+b^2}\]\[\implies \frac{a^2-b^2}{a^2+b^2}\] \[\text{Your Question is Done : )}\]

OpenStudy (anonymous):

wow...Looks so Beautiful ..\(\LaTeX\)

OpenStudy (anonymous):

Aravind !!!!

OpenStudy (aravindg):

thnk youuuuuuuuuuuuuuuuuuuuu

OpenStudy (aravindg):

ax/cos theta)+(by/sin theta)=a^2-b^2 and (axsin theta/cos^2 theta)-(bycos theta/sin^2 theta)=0 show that (ax)^2/3 +(by)^2/3=(a^2-b^2)^2/3

OpenStudy (aravindg):

can u help this one ???i got answer by taking long steps ...bt i feel there is an easier way out can u help??